How fast is 360,000 cm/hr in m/s? Use conversion factors to cancel out one unit and introduce a new unit.
Answer by Theo(13342)

i presume that cm equals centimeters and m equals meters.

you need to know the conversion factors first of all.

1 meter is equal to 10 centimeters

this means that 1 centimeters is equal to 1/10 = .1 meter.

1 hour is equal to 60 minutes
1 minute is equal to 60 seconds
1 hour is therefore equal to 60 * 60 = 3600 seconds.

this means that 1 second is equal to 1/3600 = .0002777778 of an hour.


you start with 360,000 centimeters per hour.

since 1 meters is equal to 10 centimeters, then you need to divide centimeters by 10 in order to get equivalent meters.


360,000 centimeters / 10 = 36,000 meters.

your 360,000 centimeters is equivalent to 36,000 meters.

since 1 hour is equal to 3600 seconds, you need to multiply 1 hour by 3600 to get 3600 seconds.

your 1 hour has been 3600 seconds.

360,000 centimeters per hour becomes 36,000 meters per 3600 seconds.

to find the number of meters per second you need to divide 36,000 by 3600 to get 10 meters per second.

you 360,000 centimeters per hour becomes 36,000 meters per 3600 seconds which becomes 10 meters per second.

that should be your answer.

360,000 centimeters per hour is equivalent to 10 meters per second.

to see if this is good, try to get back to centimeters per hour from this equation.

since 1 meter is equal to 10 centimeters, multiply 10 by 10 to get 100 centimeters per second.

10 meters per second becomes equivalent to 100 centimeters per second.

since there are 3600 seconds in an hour, multiply 100 centimeters per second by 3600 to get 360,000 centimeters per hour.

the conversion have been done successfully.

360,000 centimeters per hour is equivalent to 10 meters per second.